OpenStudy (anonymous):

During the first part of a trip, a canoeist travels 30 miles at a certain speed. The canoeist travels 12 miles on the second part of the trip at a speed 5 mph slower. The total time for the trip is 2 hrs. What was the speed on each part of the trip? Please round to the nearest hundredth. I am really stuck and am in need of the answers. I have worked this problem many times and keep getting the first part wrong.

OpenStudy (anonymous):

well everytime I did the math you gave me it came up wrong

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

I get the second part but not the first part. If I can get the answer to the first part I will be okay.

OpenStudy (anonymous):

It is the first part I am getting wrong.

OpenStudy (anonymous):

work back words...

OpenStudy (anonymous):

I have worked this problem so many time my head hurts. I can never get the first part of the problem right.

OpenStudy (dumbcow):

well if you know the speed of 2nd part, just add 5 to get speed of 1st part he was going 5mph slower on 2nd part

OpenStudy (anonymous):

yeah that is what is confusing to me

OpenStudy (dumbcow):

ok say he went 10mph for 1st trip then he went 5mph for 2nd trip or in general speed r in 1st trip, (r-5) in 2nd trip

OpenStudy (anonymous):

my head hurts

OpenStudy (dumbcow):

haha im sorry ok what answer did you get, we can check to see if its right

OpenStudy (anonymous):

28.46 for the first and 23.46 for the second but it tells me it is wrong

OpenStudy (dumbcow):

distance = rate*time 1st part: 30 = 28.46 * t1 t1 = 1.05 2nd part: 12 = 23.46*t2 t2 = 0.51 we know t1+t2 = 2 however 1.05+0.51 = 1.56 thats why its wrong

OpenStudy (anonymous):

man dumbcow you are amazing!!!

OpenStudy (anonymous):

okay so I am off by not much

OpenStudy (dumbcow):

haha, i've been doing this for awhile

OpenStudy (anonymous):

yeah not me. I stink at math.

OpenStudy (dumbcow):

ok so go back and replace those numbers with r and (r-5) that will represent the speeds try to solve for r again

OpenStudy (dumbcow):

the equations should be t1 = 30/r t2 = 12/(r-5) t1 +t2 = 2

OpenStudy (anonymous):

okay I feel stupid r stands for?

OpenStudy (anonymous):

sorry

OpenStudy (dumbcow):

speed of 1st trip

OpenStudy (dumbcow):

use whatever variable you want really

OpenStudy (anonymous):

okay this is the only problem I have been stuck on I will just take the hit my brain hurts and I am getting tired

OpenStudy (dumbcow):

haha ok i'll show you the solution, just try to work it out on your own later

OpenStudy (anonymous):

okay thank you.

OpenStudy (dumbcow):

Substitute in equations for t1 and t2 \[\frac{30}{r} + \frac{12}{r-5} = 2\] combine fractions, get common denominator \[\frac{30}{r}*\frac{r-5}{r-5} + \frac{r}{r}*\frac{12}{r-5} = 2\] \[\frac{30(r-5) + 12r}{r(r-5)} = 2\] Bring denominator to other side \[30(r-5) +12r =2r(r-5)\] \[30r -150 +12r = 2r^{2} -10r\] Set equal to zero \[2r^{2}-52r+150 = 0\] Factor out a 2 \[r^{2} -26r +75=0 \] Use quadratic formula, a =1, b=-26, c =75 \[r = 3.3, 22.695\] r must be greater than 5 \[r = 22.695\]

OpenStudy (anonymous):

Again thank you. I will copy this and work it over and over till I understand it.